OpenStudy (bluemiku):

Find the GCF of 56 and 40. A. 2 B. 4 C. 8 D. 280

OpenStudy (anonymous):

c.8

OpenStudy (anonymous):

sorry 4 bothering but ooooooooo a mike fan xD

OpenStudy (anonymous):

miku*

OpenStudy (anonymous):

One way to determine the greatest common factor is to find all the factors of the numbers and compare them. The factors of 40 are 1, 2, 4, 5, 8, 10, 20, and 40. The factors of 56 are 1, 2, 4, 7, 8, 14, 28, and 56. The common factors are 1, 2, 4, and 8. Therefore, the greatest common factor is 8. The greatest common factor can also be calculated by identifying the common prime factors and multiplying them together. The prime factors of 40 are 2, 2, 2, and 5. The prime factors of 56 are 2, 2, 2, and 7. The prime factors in common are 2, 2, and 2, so the greatest common factor is 2 x 2 x 2 = 8.

OpenStudy (bluemiku):

i knew it was c.!

OpenStudy (bluemiku):

me to@Survivor

OpenStudy (bluemiku):

@Survivor

OpenStudy (anonymous):

Another way to do this is with prime factorization. \[ 56 = 2^3\times 7\\ 40 = 2^3\times 5 \]So GCD is \(2^3=8\).